OpenStudy (anonymous):

What number must be subtracted from -3 to make the result zero? A) -9 B) -3 C) 3 D) 9

OpenStudy (waleed_imtiaz):

it will be +3.....

jimthompson5910 (jim_thompson5910):

no, it's -3 -3 - x = 0 -3 - x + 3 = 0 + 3 -x = +3 x = -3

jimthompson5910 (jim_thompson5910):

So -3 - (-3) = -3+3 = 0

OpenStudy (anonymous):

so its _3?

jimthompson5910 (jim_thompson5910):

-3

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ok! thanks!!!